# CamI2V: Camera-Controlled Image-to-Video Diffusion Model

### v1

@@ 5.2 Quantitative Comparison
We compare our CAMI2V with the latest methods in camera controlled image-to-video generation, including DynamiCrafter (Xing et al., 2023), MotionCtrl (Wang et al., 2024d) and CameraCtrl (He et al., 2024a). As reported in Table 1, our CAMI2V significantly improves the camera controllability, with substantial reductions in RotErr, TransErr and CamMC. Compared to MotionCtrl, our method reduces RotErr by 0.512, translating to a 3.67 decrease in rotational error, which marks a significant improvement. However, enhancing camera controllability inevitably leads to a decline in FVD. Nevertheless, our method surpasses the state-of-the-art method CameraCtrl in both camera controllability and FVD metrics.
@@ 5.3 Ablation Study
Table 1: Quantitative comparison with state-of-the-art methods. Our method achieves a significant improvement in camera controllability (25.5% in rotation and 7.77% in translation compared to previous sota), where a decrease of 0.1 in RotErr corresponds to a reduction of 0.72 degrees in rotational error. * denotes the results we reproduced using DynamiCrafter as base I2V model.
@@
DynamiCrafter ECCV 2024 $9.674$ $3.086$ $11.28$ $113.25$ $97.181$
DynamiCrafter+MotionCtrl* SIGGRAPH 2024 $2.632$ $0.955$ $3.197$ $\bf 72.073$ $\bf 62.069$
DynamiCrafter+CameraCtrl* arXiv 2024 $1.636$ $0.595$ $1.958$ $78.795$ $66.916$
DynamiCrafter+CamI2V (ours) - $\bf 1.509$ $\bf 0.443$ $\bf 1.726$ $77.265$ $65.806$
Table 2: Ablation study on model variants.
@@
VideoGPT StyleGAN
DynamiCrafter+CamI2V (ours) $\bf 1.509$ $\bf 0.443$ $\bf 1.726$ $77.265$ $65.806$
$1.605$ $0.497$ $1.851$ $\bf 70.068$ $\bf 60.830$
$1.516$ $0.482$ $1.757$ $72.069$ $61.999$
$4.613$ $1.844$ $5.799$ $107.97$ $91.318$
$6.627$ $3.677$ $9.287$ $115.36$ $100.58$
DynamiCrafter $10.10$ $3.268$ $11.79$ $127.41$ $107.12$
We conduct ablation study of proposed method on Plücker embedding, 3D full attention and our proposed epipolar attention. All model variants are trained under the same protocol mentioned in Section 5.1 for 50K steps. Evaluation results under CFG $s_{\text{img\&txt}}=4$ and $s_{\text{camera}}=1.875$ are provided in Table 2. In this section, we will discuss the validity of our Plücker pose encoder and epipolar-constrained attention module.
Plücker embedding. The experimental results illustrate that Plücker embedding significantly enhance the capability of base model for camera control. It also coherently improve visual quality of generated frames, witnessed by a large drop of FVD. We conclude that Plücker embedding serves as a more expressive position encoding and is more effective for the model to leverage geometric correspondence, thus better cross-frame consistency.
3D full attention alone without Plücker embedding. The use of 3D Full Attention alone, without employing Plücker embeddings to provide global positional encoding (a form of clean condition), fails to achieve camera controllability. This indicates that even with a 3D global receptive field, the presence of noise limits the model’s ability to generate videos that follow specific camera trajectories, resulting instead in random video generation that conforms to the dataset distribution.
Epipolar attention alone without Plücker embedding. The use of Epipolar Attention alone, without Plücker embeddings for global positional encoding (a form of clean condition), still retains a degree of camera controllability. This notable phenomenon suggests that explicit constraints can enable the model to grasp knowledge about the physical world.
3D full attention along with Plücker embedding. The use of 3D full attention with employing Plücker embeddings to provide global positional encoding (a form of clean condition), achieves the best FVD but results in the poorest camera controllability metrics. This supports our earlier discussion in the introduction that 3D full attention loses its ability to track features under high noise conditions. However, in small noise scenarios, it can access a greater number of features, allowing for the generation of videos that conform to the dataset distribution, though these may not align with the camera trajectories provided by users.
Epipolar attention along with Plücker embedding. When the explicit constraint of Epipolar Attention is combined with the implicit prior provided by Plücker embeddings for global positional encoding, the best camera controllability is achieved. This underscores the necessity of explicit constraints like Epipolar Attention. Although the FVD declines, this is understandable; when FVD reaches a saturation point, enhancing controllability inevitably leads to deviations from the current distribution. Achieving the best FVD with 3D Full Attention alone, without controllability, lacks practical value in the field of controllable generation.
@@ Appendix C GPU Memory and Speed
Table 3: Comparison on GPU memory usage and speed under DeepSpeed ZeRO-1. * denotes our reproduction on DynamiCrafter. We report full parameter fine-tuning results of DynamiCrafter. Our model can be trained on 24GB custom-level GPUs despite the additional epipolar attention.
Method # Trainable Parameters GPU Memory (GiB) $\downarrow$ Time (s) $\downarrow$
@@ Appendix D Extra Out-of-Domain Visualizations
Figure 9: Orbit right on unseen images.
Figure 10: Zoom in on unseen images.
Figure 11: Orbit left on unseen images.
Figure 12: Zoom out on unseen images.
Appendix E Extra RealEstate10K Visualizations
Figure 13: Extra RealEstate10K Visualizations.

### v3

@@ 5 Experiments
Table 1: Quantitative comparison with state-of-the-art methods. * denotes the results we reproduced using DynamiCrafter as base I2V model. We achieve a 32.96%, 25.64%, 20.77% improvement over previous Sota CameraCtrl on RotErr, CamMC, TransErr on the RealEstate10K dataset without compromising dynamics, generation quality, and generalization on out-of-domain images. These results were obtained using Text and Image CFG set to 7.5, 25 steps, and camera CFG set to 1.0 (no camera cfg).
Method Publication TransErr $\downarrow$ RotErr $\downarrow$ CamMC $\downarrow$ FVD $\downarrow$
VideoGPT StyleGAN
DynamiCrafter (Xing et al., 2023) ECCV 2024 9.8024 3.3415 11.625 106.02 92.196
+ MotionCtrl (Wang et al., 2024d)* SIGGRAPH 2024 2.5068 0.8636 2.9536 70.820 60.363
+ CameraCtrl (He et al., 2024a)* arXiv 2024 1.9379 0.7064 2.3070 66.713 57.644
+ CamI2V (Ours) 1.4955 0.4758 1.7153 66.090 55.701
@@ 5.2 Quantitative Comparison
We compare our CamI2V with the latest methods in camera controlled image-to-video generation, including DynamiCrafter (Xing et al., 2023), MotionCtrl (Wang et al., 2024d) and CameraCtrl (He et al., 2024a). As reported in Table 1, our CamI2V significantly improves the camera controllability and visual quality, with substantial reductions in RotErr, TransErr, CamMC and FVD. Compared to CameraCtrl, our method reduces RotErr by $0.2306$, translating to a $13.21^{\circ}$ decrease in rotational error, which marks a significant improvement. And our method surpasses the state-of-the-art method CameraCtrl in other camera controllability and FVD metrics.
@@ 5.3 Ablation Study
Method Plücker Epipolar 3D Full TransErr $\downarrow$ RotErr $\downarrow$ CamMC $\downarrow$ FVD $\downarrow$
@@
DynamiCrafter +CamI2V (Ours) 1.4955 0.4758 1.7153 66.090 55.701
@@
1.8215 0.6299 2.1315 71.026 60.000
1.8877 0.7098 2.2557 66.077 55.889
5.5119 1.3988 6.2855 92.605 81.447
DynamiCrafter 9.8024 3.3415 11.625 106.02 92.196
Adding more conditions to generative models typically reduces uncertainty and improves generation quality (e.g. providing detailed text conditions through recaption). In this paper, we argue that it is also crucial to consider noisy conditions like latent features $z_{t}$, which contain valuable information along with random noise. For instance, in SDEdit (Meng et al., 2021) for image-to-image translation, random noise is added to the input $z_{0}$ to produce a noisy $z_{t}$. The clean component $z_{0}$ preserves overall similarity, while the introduced noise leads to uncertainty, enabling diverse and varied generations.
In this paper, we argue that providing the model with more noisy conditions, especially at high noise levels, does not necessarily reduce more uncertainty, as the noise also introduces randomness and misleadingness. This is the key insight we aim to convey.
To validate this point, we designed experiment with the following setups:
  1. 1.
    Plücker Embedding (Baseline): This setup, akin to CameraCtrl, has minimal noisy conditions on cross frames due to the inefficiency of the indirect cross-frame interaction (spatial and temporal attention).
  2. 2.
    Plücker Embedding + Epipolar Attention only on reference frame: Similar to CamCo, this setup treats the reference frame as the source view, enabling the target frame to refer to it. It accesses a small amount of noisy conditions on the reference frame. However, some pixels of the current frame may have no epipolar line interacted with reference frame, causing it to degenerate to a CameraCtrl-like model without epipolar attention.
  3. 3.
    Plücker Embedding + Epipolar Attention (Our CamI2V): This setup can impose epipolar constraints with all frames, including adjacent frames that have interactions in most cases to ensure an sufficient amount of noisy conditions.
  4. 4.
    Plücker Embedding + 3D Full Attention: This configuration allows the model to directly interact with features of all other frames, accessing the most noisy conditions.
The amount of accessible noisy conditions of the above four setups increase progressively. One might expect that 3D full attention, which accesses the most noisy conditions, would achieve the best performance. However, as shown in Tab. 2, 3D full attention performs only slightly better than CameraCtrl and is inferior to CamCo-like setup who only applies epipolar attention on reference frame. Notably, our method achieves best result by interacting with more noisy conditions along the epipolar lines. It can be clearly seen in the comparison part in supplementary that CamCo-like setup reference much on the first frame and cannot generate new objects. The 3D full attention generates objects within large movement due to its access to all frames pixels while it is affected by incorrect position of pixels. These findings confirm our insight that an optimal amount of noisy conditions leads to better uncertainty reduction, rather than merely increasing the quantity of noisy conditions.
@@ Appendix C GPU Memory and Speed
Table 3: Comparison on GPU memory usage and speed under DeepSpeed ZeRO-1. * denotes our reproduction on DynamiCrafter. We report full parameter fine-tuning results of DynamiCrafter. Our model can be trained on 24GB consumer-level GPUs despite the additional epipolar attention.
Method # Params Trainable GPU Memory (GiB) $\downarrow$ Time (s) $\downarrow$
@@ Appendix D Extra Out-of-Domain Visualizations
Dynamic videos are best viewed at our local anonymous web page. It’s strongly recommended to view the visualizations in the supplementary for a more comprehensive evaluation.
Figure 9: Visualization of our 256$\times$256 model.
Figure 10: Visualization of original outputs from our 512$\times$320 model, with no padding removed.
Figure 11: Generated by our 512$\times$320 model, compatible with input images of arbitary aspect ratio.